# Anomalous Change Point Detection Using Probabilistic Predictive Coding

### v1

@@ Abstract
Change point detection (CPD) and anomaly detection (AD) are essential techniques in various fields to identify abrupt changes or abnormal data instances. However, existing methods are often constrained to univariate data, face scalability challenges with large datasets due to computational demands, and experience reduced performance with high-dimensional or intricate data, as well as hidden anomalies. Furthermore, they often lack interpretability and adaptability to domain-specific knowledge, which limits their versatility across different fields. In this work, we propose a deep learning-based CPD/AD method called Probabilistic Predictive Coding (PPC) that jointly learns to encode sequential data to low dimensional latent space representations and to predict the subsequent data representations as well as the corresponding prediction uncertainties. The model parameters are optimized with maximum likelihood estimation by comparing these predictions with the true encodings. At the time of application, the true and predicted encodings are used to determine the probability of conformity, an interpretable and meaningful anomaly score. Furthermore, our approach has linear time complexity, scalability issues are prevented, and the method can easily be adjusted to a wide range of data types and intricate applications. We demonstrate the effectiveness and adaptability of our proposed method across synthetic time series experiments, image data, and real-world magnetic resonance spectroscopic imaging data.
@@ 1 Introduction
Change point detection (CPD) and anomaly detection (AD) are critical techniques in various fields, involving the detection of abrupt changes or abnormal instances in sequential data, respectively. Both techniques are widely used in finance surveillance and fraud detection [21], Internet of Things and (remote) health monitoring [13], video surveillance [11], and intrusion detection in cyber security systems [66]. With ever-increasing amounts of data, these techniques have become more important, as they decrease labor efforts and enable the detection of subtle but significant anomalies and change points.
@@
In this work, we propose Probabilistic Predictive Coding (PPC), a method that does not suffer from these limitations. PPC is a deep learning-based CPD/AD method that encodes sequential data to latent space representations and thus can be adapted to a wide range of data types, including time series, images, videos, graphs, sets, and text data. Due to the application of deep learning, the computational complexity at application time is linear to the number of samples, preventing scalability issues. At the same time, the deep learning approach makes the method widely applicable in different domains with different levels of data complexity, while still being able to be tuned to domain knowledge. By using predictability modeling, the method is intuitive and interpretable. Furthermore, due to its probabilistic nature, the method can provide a meaningful anomaly score in the form of a probability of conformance.
@@ 2.1 Applications of CPD and AD
2.2 Review of existing methodologies
Most surveys on anomaly detection make a clear distinction between conventional/classical/traditional methods and deep learning-based methods. As identified by [46], previous surveys tend to focus on conventional methods, while the research field of deep anomaly detection is very active. In this section, we will provide a short survey of existing classical and deep learning-based methods, including relevant strengths and weaknesses.
2.2.1 Classical methods
Recently, [55] surveyed traditional anomaly detection methods, where they sorted these algorithms into several categories. We refer the reader to this work for extensive information on this topic but will summarize the categories.
Statistical models like Grubbs’s test [18] or $\chi^{2}$-statistics [68] use statistical tests to find anomalies and provide statistically meaningful scores. Non-parametric statistical methods like histogram-based methods, e.g., [15], and kernel-based algorithms such as proposed by [35], both apply a form of density estimation, assuming that anomalies have low density. While these methods are simple and intuitive, they have difficulties processing highly complex data.
Another density-based method is Local Outlier Factor (LOF) [6], which calculates the relative density of data instances with respect to their neighboring data instances. A high number of variants on LOF have been proposed, such as Resolution-based Outlier Factor (ROF) [14], Global-Local Outlier Scores from Hierarchies (GLOSH) [7] and Simple univariate Probabilistic Anomaly Detector (SPAD) [3]. While these methods generally perform well, they are computationally expensive and are not suitable for large or high-dimensional datasets.
Distance-based methods assume that anomalous data instances lie far away from their nearest neighbors. Well-known examples are the k-nearest neighbors algorithm [27] and AntiHub [53]. These methods are relatively easy to implement but are computationally expensive, less effective in large or high-dimensional datasets, and do not provide intuitive anomaly scores.
Clustering-based methods assume that anomalous data instances are not part of a cluster of data instances. By clustering a collection of data instances, outliers can be detected. Examples of methods are OFP (Outlier Finding Process) [25] and Clustering-based Outlier Detection (CBOD) [26]. These methods are adaptable to different data types and intricate problem setups. However, the performance of these types of methods is generally sensitive to the type of clustering algorithm and hyperparameters used, and in general, clustering is computationally expensive.
Isolation-based algorithms assume that anomalous data instances are easier to isolate than normal data instances when splitting data into partitions. Common methods are isolation forests [36] and Unsupervised Stochastic Forest-based Anomaly Detector (usfAD) [2]. While these methods have relatively low computational complexity compared to clustering- or distance-based methods, they do not provide meaningful or intuitive anomaly scores, as the scores lack context and are not normalized or standardized.
Lastly, subspace-based methods assume that anomalous data instances are hidden in a subset of features. By selecting a subspace and computing a score for that subspace, anomalies may be detected. Examples of such methods are Subspace Outlier Degree (SOD) [30] and Zero++ [48]. While these methods are generally well able to detect anomalies concealed in high-dimensional spaces, they have high computational costs, similar to distance- and density-based methods.
2.2.2 Deep learning-based methods
In the recent work of [46], the authors proposed a taxonomy for deep anomaly detection techniques, consisting of three main categories of techniques. The main difference between these categories is the degree to which the anomaly score is part of the model training pipeline. We refer the reader to this work for more details but will summarize this taxonomy and give the advantages and disadvantages of current approaches.
Whenever the anomaly scoring mechanism and the deep learning model are independently established, the authors refer to deep learning for feature extraction. The main objective is to apply nonlinear dimensionality reduction instead of more traditional linear methods like principal component analysis (PCA). In that sense, deep learning techniques have been shown to extract significantly richer features [16], which can subsequently be used with any of the classical methods mentioned in section 2.2.1. While these methods are generally easy to implement and pre-trained deep learning models are widely available, disjoint development of the feature extraction and anomaly scoring mechanisms often leads to sub-optimal performance [46].
On the other end of the spectrum, end-to-end anomaly score training aims to fully include the anomaly score in the training pipeline. No existing or pre-defined anomaly measures are needed with this approach, as the network itself has to learn the anomaly scores directly. Examples within this category include ranking models (e.g., LeSiNN, [47]), prior-driven models (e.g., [43]), softmax likelihood models (e.g., [8]), and end-to-end one-class classification methods (e.g., [54]). One major advantage of these methods is that direct optimization of the anomaly scores is possible, often leading to great performance. However, they generally require weak supervision and some form of labeled anomalies. Furthermore, providing these labeled anomalies may inhibit generalization to unseen anomalies.
The last proposed category is somewhere in between the previous two categories. It aims to learn feature representations of non-anomalous data instances, hence the name learning feature representations of normality. Some examples of generic methods that capture data regularities are autoencoders (e.g., variational autoencoders, [10]), generative adversarial networks (e.g., f-AnoGAN [56]), predictability modeling (e.g., AnoPCN [67]), and self-supervised classification methods (e.g., [60]). Additionally, some methods are specifically tuned to certain anomaly measures, such as distance-based measures (e.g., [63]), one-class classification measures (e.g., [64]), and clustering-based measures (e.g., [65]). Most of these methods are easy to implement and only require semi-supervised training with non-anomalous data. However, these methods generally lack clear interpretability, provide anomaly scores without sufficient contextual information or a standardized scale, or are computationally complex and expensive.
Closest to our proposed method is the VAE-LSTM method introduced by [34], falling in the category learning feature representations of normality. VAE-LSTM uses a hybrid method consisting of variational autoencoders (VAE) and long short-term memory (LSTM) to detect anomalies. Similar to PPC, it aims to encode data to latent space representations and predict future representations using the LSTM. However, the VAE and LSTM models are trained disjointly, which yields latent space representations that are not optimized for prediction purposes [46]. In our method, we jointly train the encoder model and prediction models to enhance the anomaly detection performance. Furthermore, VAE-LSTM uses a decoder to reconstruct the original data instances based on the (predicted) latent space representations, where the reconstruction error is used as the anomaly score. We opted to calculate a more meaningful and intuitive anomaly score by obtaining probabilistic predictions of latent space representations and comparing them with the true representations, yielding a probability of conformance.
@@ 3.1 Problem definition
Let $\{x_{1},x_{2},\cdots\}$ be a set consisting of data instances of any type that are sequential, predictable, and encodable. Data instances are sequential whenever they are part of a unidirectional succession, e.g., $x_{2}$ succeeds $x_{1}$, $x_{3}$ succeeds $x_{2}$, etc. The property of predictability is applicable whenever individual data instances can be reasonably estimated given prior data instances in the sequence, indicated by a relatively high entropy of the conditional probability density function $f\left(x_{j+i}|x_{1},x_{2},\cdots,x_{j}\right)$ for prediction step $i$. Data instances are encodable whenever all data instances $\{x_{1},x_{2},\cdots\}$ can be transformed to low-dimensional latent space representations $\{z_{1},z_{2},\cdots\}$ with an encoder neural network, and these latent space representations contain the necessary information to satisfy the requirement of predictability. Obvious examples of data that meet these requirements are video (e.g., with convolutional neural networks) or text (e.g., with Word2vec), but certain time series, images, and even graphs may qualify as well. It is up to the user to define a sequential axis within the data such that the data instances indeed have these properties.
@@ 3.2 Probabilistic predictions
Similar to the CPC pipeline, our pipeline uses an encoder neural network $E$ to encode the data instances into low-dimensional latent space encodings, i.e., $z_{k}=E\left(x_{k}\right)$. A recurrent neural network $G$ translates multiple sequential encodings to a context vector $c_{N_{p}}$, noted as $c_{N_{p}}=G\left(z_{1},z_{2},\cdots,z_{N_{p}}\right)$. This context vector then forms the basis for predictions of future latent space representations, where several forecasting neural networks $F_{1},F_{2},\cdots,F_{N_{f}}$ transform this context vector into predictions for the next $N_{f}$ latent space representations, i.e., $\left(\hat{z}_{N_{p}+i},\sigma_{N_{p}+i}\right)=F_{i}\left(c_{N_{p}}\right)$.
While CPC uses single-layer perceptrons without estimating prediction uncertainty, we opted to use mean-variance estimation neural networks that employ multi-task learning to estimate both the encoding $\hat{z}_{N_{p}+i}$ and the associated prediction uncertainty $\sigma_{N_{p}+i}$ [42]. We opted with the proposed PPC method for probabilistic predictions because (1) this allows for calculating a proxy of the likelihood $f\left(x_{N_{p}+i}|x_{1},x_{2},\cdots,x_{N_{p}}\right)$ (see section 3.3), (2) this likelihood is an intuitive and non-arbitrary distance metric between ground truth and predicted encodings and forms a suitable loss function for training neural networks (see section 3.4), and (3) probabilistic predictions allow for a nuanced and statistically meaningful interpretation of the output in the form of a probability of conformance (see section 3.5). Figure 1 shows an example of the PPC pipeline architecture.

### v2

@@ Abstract
Change point detection (CPD) and anomaly detection (AD) are essential techniques in various fields to identify abrupt changes or abnormal data instances. However, existing methods are often constrained to univariate data, face scalability challenges with large datasets due to computational demands, and experience reduced performance with high-dimensional or intricate data, as well as hidden anomalies. Furthermore, they often lack interpretability and adaptability to domain-specific knowledge, which limits their versatility across different fields. In this work, we propose a deep learning-based CPD/AD method called Probabilistic Predictive Coding (PPC) that jointly learns to encode sequential data to low-dimensional latent space representations and to predict the subsequent data representations as well as the corresponding prediction uncertainties. The model parameters are optimized with maximum likelihood estimation by comparing these predictions with the true encodings. At the time of application, the true and predicted encodings are used to determine the probability of conformance, an interpretable and meaningful anomaly score. Furthermore, our approach has linear time complexity, scalability issues are prevented, and the method can easily be adjusted to a wide range of data types and intricate applications. We demonstrate the effectiveness and adaptability of our proposed method across synthetic time series experiments, image data, and real-world magnetic resonance spectroscopic imaging data.
@@ 1 Introduction
Change point detection (CPD) and anomaly detection (AD) are critical techniques in various fields, involving the detection of abrupt changes or abnormal instances in sequential data, respectively. Both techniques are widely used in financial surveillance and fraud detection [31], Internet of Things and (remote) health monitoring [21], video surveillance [19], and intrusion detection in cybersecurity systems [90]. With ever-increasing amounts of data, these techniques have become more important, as they decrease labor efforts and enable the detection of subtle but significant anomalies and change points.
@@
In this work, we propose Probabilistic Predictive Coding (PPC), a method that does not suffer from these limitations. PPC is a deep learning-based CPD/AD method that encodes sequential data to latent space representations and thus can be adapted to a wide range of data types, including time series, images, videos, graphs, sets, and text data. Due to the application of deep learning, the computational complexity at application time is linear with respect to the number of samples, preventing scalability issues. At the same time, the deep learning approach makes the method widely applicable in different domains with different levels of data complexity, while still being able to be tuned to domain knowledge. By using predictability modeling, the method is intuitive and interpretable. Furthermore, due to its probabilistic nature, the method can provide a meaningful anomaly score in the form of a probability of conformance.
@@ 2.1 Applications of CPD and AD
2.2 Anomaly detection methods
Anomaly detection methods aim to identify data instances that deviate significantly from the expected behavior of normal data. Over the decades, approaches have evolved from statistical hypothesis tests to probabilistic and deep learning–based methods, reflecting a shift from static modeling to dynamic, data-driven prediction. While we will mention some highlights of the developments over the years, we refer the reader to comprehensive reviews like the work of [66] for in-depth information.
Early statistical approaches relied on parametric hypothesis testing and robust statistics. Grubbs’ test [27] detects univariate outliers via Studentized range thresholds, while the multivariate $\chi^{2}$-AD method [91] generalizes this principle using chi-squared statistics for multivariate or network data. Such models offer interpretability and computational efficiency but depend on assumptions of normality and independence that rarely hold in complex, non-stationary settings. Early time-series models, such as ARIMA [10] and Exponential Smoothing (ES, DES, TES; 75), extended this paradigm to temporal prediction, but remained limited to linear dynamics and Gaussian noise.
Proximity- and density-based methods reframed anomalies as geometrically or statistically isolated points in feature spaces. Distance-Based Outlier Detection (DBOD, 38) identifies points with sparse $R$-neighbourhoods, while Local Outlier Factor (LOF, 11) measures local reachability density relative to neighbors. Histogram-Based Outlier Detection (HBOD, 24) and HDBSCAN [12] improve scalability and hierarchical robustness.
Clustering-based methods such as Two-Phase Clustering (TPC, 35) and Clustering-Based Outlier Detection (CBOD, 36) interpret anomalies as data points that do not belong to any dense cluster. While effective for low-dimensional, static data, these methods are ineffective in high-dimensional or evolving distributions.
Subspace and ensemble approaches improved scalability and robustness. The Subspace Outlier Degree (SOD, 41) identifies anomalies within relevant feature subspaces, while ensemble frameworks such as ZERO++ [65] combine feature projections for stability. Isolation Forest (iForest, 47) isolates anomalies by randomly partitioning data. However, these techniques generally ignore temporal dependencies.
Forecasting and reservoir-based sequence models emerged to handle sequential data. Echo State Networks (CoalESN, MoteESN; 60, 13) applied recurrent reservoirs for streaming anomaly detection, while Probabilistic Change Identification (PCI, 92) incorporated adaptive thresholding. Yet, these methods often assume stationarity and lack principled uncertainty estimation.
Deep representation learning introduced end-to-end feature extraction. Recurrent architectures such as LSTM-AD [51] and EncDec-AD [50] train encoder–decoder LSTMs on normal sequences, detecting anomalies via reconstruction error. OmniAnomaly [76] extends this with stochastic latent variables for robustness.
Generative probabilistic models such as VAE-LSTM [46] model normal dynamics through probabilistic reconstruction, while GAN-based variants like f-AnoGAN [72], TadGAN [25], and TAnoGAN [6] learn adversarial latent representations of normality. Hybrids such as USAD [4] jointly optimize reconstruction and prediction losses to tighten anomaly bounds, though thresholding remains heuristic.
Transformer-based methods have recently achieved state-of-the-art performance on multivariate and long-range dependencies. MTAD-GAT [95] incorporates graph attention for inter-variable relations, Anomaly Transformer [88] quantifies associations via self-attention discrepancy, and TranAD [81] and PatchAD [96] use efficient patch-level encoders for scalable reconstruction. TimesNet [87] reformulates 1D series as 2D temporal patterns, while AnomalyLLM [49] fine-tunes large language models for zero-shot anomaly reasoning.
2.3 Change point detection methods
Change point detection (CPD) aims to identify points in time where the properties of a process change abruptly, causing significant shifts in sequential data. Like AD methods, CPD methods have evolved from simpler statistical tests to deep learning-based solutions. Highlights of developments are mentioned here, but more comprehensive surveys can be found in [80] and [89].
Early CPD research focused on sequential likelihood-ratio methods. The cumulative sum (CUSUM, 63) test introduces recursive accumulation of log-likelihood differences between competing hypotheses of stationarity and change. Together with related sequential probability ratio tests [83], CUSUM forms the foundation of online detection theory, defining optimality in terms of the average run length to false alarm.
Later developments reinterpreted such sequential schemes through a Bayesian lens. The Bayesian Online Changepoint Detection (BOCPD, 1) algorithm expresses the posterior over the run length since the last change as a recursive message-passing update weighted by the model’s predictive likelihood. BOCPD established the modern probabilistic framing of online CPD and inspired extensions to streaming, high-dimensional, and budgeted settings, e.g., [84].
Parallel to sequential approaches, research was focused on offline change point detection, where the entire data sequence is observed and optimal segmentation is desired. Early econometric contributions, such as [5], formulated CPD as a model-selection problem over linear regression segments, combining Wald-type test statistics with trimming strategies to ensure asymptotic consistency.
Subsequent research developed exact search algorithms under penalized cost formulations. The Optimal Partitioning and Pruned Exact Linear Time (PELT, 37) method achieves linear computational cost by pruning candidate boundaries while maintaining global optimality under additive penalties. Randomized algorithms such as Wild Binary Segmentation (WBS; 23) improved practical sensitivity to localized or short-lived changes by evaluating CUSUM statistics over random subintervals rather than fixed partitions. Together, these methods established the dominant penalized optimization paradigm for multiple change point inference.
While parametric likelihood formulations dominate classical CPD, many real-world applications exhibit changes beyond shifts in mean or variance. Nonparametric methods address this by comparing empirical distributions across adjacent windows without assuming a specific distribution. The energy-distance approach of [52] detects distributional changes by maximizing between-segment dissimilarity in terms of pairwise distances, yielding consistency for multivariate and heavy-tailed data. Kernel-based detectors, notably those by [29], leverage reproducing kernel Hilbert space embeddings to detect nonlinear and high-order distributional changes via maximum mean discrepancy or Hilbert–Schmidt independence criteria. These approaches generalize the notion of a contrast function to measure divergence between local empirical distributions.
Recent research has expanded CPD to structured and high-dimensional domains. The Group Fused Lasso [8] enforces joint sparsity across multiple correlated series, identifying shared change points via total-variation regularization. High-dimensional theoretical analyses, such as those by [15], derive consistency guarantees for sparse change recovery in regimes where the number of variables exceeds the sample size. Sliding-window methods like MOSUM [16] and its multivariate extensions, e.g., [55], adapt moving-sum test statistics to high-dimensional covariance structures, allowing scalable online operation. These approaches extend classical univariate tests to complex multivariate and structured data.
In parallel, modern representation-learning techniques employ learned or neural approaches to CPD. Deep architectures for time series segmentation now frame change detection as an end-to-end learning problem, where neural networks infer latent representations that linearize temporal regimes and produce differentiable change scores [45, 58]. Other recent work explores continuous-time latent stochastic differential equation models for CPD [71], merging neural dynamics with Bayesian inference. Although many deep methods achieve strong performance, they typically depend on supervised labels or synthetic data generation to train segmentation boundaries.
Recent advances have also introduced unsupervised deep learning approaches for CPD, which reduce reliance on labeled change points. Adaptive LSTM–autoencoder architectures have been proposed for memory-free online CPD [3], while graph neural network encoder–decoder models capture evolving correlations in multivariate sequences [93]. Furthermore, fully unsupervised multivariate CPD using minimum-description-length guided greedy segmentation has been developed [85]. These unsupervised methods highlight the growing emphasis on flexible, label-free detection of structural shifts in complex time series.
2.4 Anomalous change point detection methods
Anomalous change point detection (ACPD) methods not only aim to identify distributional breaks but also to determine which of those breaks are truly anomalous instead of routine regime shifts. Early works established this two-score approach. ChangeFinder [78] produces both outlier and change point scores within an online, likelihood-based framework, making a clear distinction between transient atypical observations and persistent distributional changes. Bayesian latent-variable models extend this idea by inferring posterior indicators for both event types. For example, ABCO (Adaptive Bayesian Changepoint & Outlier scoring) places local outlier processes inside a state-space forecasting model and yields posterior probabilities for changepoints and outliers [86], while recent online Bayesian models explicitly aim to jointly detect collective anomalies and change points in streaming, detailed series [14].
Other lines of work detect anomalous change points via segmentation and practical pipelines. Penalized-cost segmentation methods that separate collective (segment) anomalies from point anomalies (e.g., the linear-time CAPA family and related algorithms) provide efficient offline labelling of anomalous segments versus point events [22]. In applied monitoring contexts, pipelines that use CPD outputs to suppress anomaly alarms show strong empirical value. Apostol et al. propose a CPD-enhanced anomaly pipeline for IoT data that reduces false positives by treating change points as expected regime shifts unless further evidence indicates anomaly [2].
Recent work moves toward multivariate, non-linear, and decision-theoretic formulations. Liu et al. developed an unsupervised approach for simultaneous anomaly and change point detection under concept drift by using fluctuation-based features and persistence testing to distinguish transient anomalies from regime shifts in multivariate series [48]. On the theoretical side, formulations that seek an anomalous change in one process among many provide asymptotic optimality results for sequential search strategies [18].
These methods show clear trade-offs: classical/online methods offer real-time processing but limited representation power [78]; Bayesian models provide calibrated event posteriors but have historically assumed simple dynamics or operated offline, though recent extensions enable online processing [86, 14]; segmentation and pipeline methods scale well in practice but are often heuristic and non-probabilistic [22, 2]; and recent approaches begin to handle multivariate, drifting data but rarely combine deep representation learning with calibrated, change-point-specific anomaly scoring [48, 18].
Probabilistic Predictive Coding addresses this gap by combining learned predictive latent representations with explicit probabilistic conformance scoring for change points: it retains probabilistic modeling while providing the representation power and (online/offline) applicability needed for modern multivariate sensors, thereby enabling distinction between anomalous and routine change points.
@@ 3.1 Problem definition
Let $\{x_{1},x_{2},\cdots\}$ be a set consisting of data instances of any type that are sequential, predictable, and encodable. Data instances are sequential whenever they are part of a unidirectional succession, e.g., $x_{2}$ succeeds $x_{1}$, $x_{3}$ succeeds $x_{2}$, etc. The property of predictability is applicable whenever individual data instances can be reasonably estimated given prior data instances in the sequence, indicated by a relatively low entropy of the conditional probability density function $f\left(x_{j+i}|x_{1},x_{2},\cdots,x_{j}\right)$ for prediction step $i$. Data instances are encodable whenever all data instances $\{x_{1},x_{2},\cdots\}$ can be transformed to low-dimensional latent space representations $\{z_{1},z_{2},\cdots\}$ with an encoder neural network, and these latent space representations contain the necessary information to satisfy the requirement of predictability. Obvious examples of data that meet these requirements are video (e.g., with convolutional neural networks) or text (e.g., with Word2vec), but certain time series, images, and even graphs may qualify as well. It is up to the user to define a sequential axis within the data such that the data instances indeed have these properties.
@@ 3.2 Probabilistic predictions
An overview of the proposed PPC pipeline is shown in Figure 1. The architecture consists of the following key modules. The model choice for each module is flexible and can be replaced with any suitable alternative.
  • Encoder ($E$): A neural network that transforms each input data instance $x_{k}$ into a low-dimensional latent representation $z_{k}=E(x_{k})$. This step compresses the input while preserving relevant features for downstream prediction. The encoder is typically implemented as a convolutional or feedforward network, depending on the data modality.
  • Sequence Model ($G$): A recurrent neural network (e.g., GRU or LSTM) that processes the sequence of latent representations $\{z_{1},z_{2},\cdots,z_{N_{p}}\}$ and summarizes them into a context vector $c_{N_{p}}=G(z_{1},z_{2},\cdots,z_{N_{p}})$. This context vector captures temporal dependencies and serves as a compact summary of the observed sequence.
[…]
  • Decoder ($D$): This module plays a vital role during the training phase but is not used during the anomaly detection inference phase. Its function is to reconstruct the original input data from its latent space representation. This reconstruction acts as a regularisation mechanism during joint training, compelling the Encoder to learn meaningful and invertible latent representations.
We opted for this probabilistic prediction framework because: (1) this allows for calculating a proxy of the likelihood $f\left(x_{N_{p}+i}|x_{1},x_{2},\cdots,x_{N_{p}}\right)$ (see section 3.3), (2) this likelihood is an intuitive and non-arbitrary distance metric between ground truth and predicted encodings and forms a suitable loss function for training neural networks (see section 3.4), and (3) probabilistic predictions allow for a nuanced and statistically meaningful interpretation of the output in the form of a probability of conformance (see section 3.5). Figure 1 shows an example of the PPC pipeline architecture.